OpenStudy (anonymous):

on an auto mechanic aptitude test the mean score is 270points and the standard deviation is 35points. is a student has standardized score of 1.9, how many points is that? if a student has a standardized score of -0.25, how many points us that? Assuming the scores follow a normal distribution, what is the probability that a student will get between 200 and 340 points?

OpenStudy (anonymous):

1.9 = (x - 270)/35 x = 336.5 b) The same as before c) P((200 - 270)/35 ≤ (X - 270)/35 ≤ (340 - 270)/35) = P( - 2 ≤ Z ≤ 2) = P(Z ≤ 2) - P(Z ≤ - 2) = 0.9772 - 0.0228 = 0.9544